OpenStudy (anonymous):

Write an equation for a circle with center at (-2,7) and diameter of 14.

OpenStudy (anonymous):

Equation of a circle \[\Large (x-h)^2+(y-k)^2=r^2\] (h,k) center r radius Simply plug in

OpenStudy (anonymous):

I still don't understand how to

OpenStudy (anonymous):

it doesn't really make since

OpenStudy (anonymous):

Well a circle's equation is that up there^, ok 

OpenStudy (anonymous):

OK see the center right there? Its the point (1,3) and we know our radius is 3 becasue we made it that way From the equation we just plug it in \[(x-(1))^2+(y-3)^2=(3)^2=~~~~~(x-1)^2+(y-3)^2=9\]

OpenStudy (anonymous):

And as you know Diameter= 2 times the radius so the radius is half of 14

OpenStudy (anonymous):

ah ok thanks

OpenStudy (anonymous):

Remember if the center is (-1,-4) then its \[\Large (x-(-4))^2=(x+4)^2\]

OpenStudy (anonymous):

I just meant if one of the coordinates are negative